based on Newton's Law of Gravitation,the force,F,between two points masses,M and m,separate by a distance r is given by

f=g multiply by Mm divided by r square
where Gis the gravttational constant

What is the base unit for F square?
1
Expert's answer
2019-07-25T17:10:12-0400
"F=\frac{GmM}{r^2}"

"[F]=\frac{[G][m][M]}{[r]^2}=\frac{(m^3kg^{-1}s^{-2})kg^2}{m^2}=m^1kg^1s^{-2}"

"[F^2]=[F]^2=m^2kg^2s^{-4}"